OpenStudy (anonymous):

add 1/14+11/12= then simplify

OpenStudy (anonymous):

\[\frac{1}{14}+\frac{11}{12}\] \[\frac{12}{168}+\frac{154}{168}\]

OpenStudy (anonymous):

\[\frac{166}{168}=\]

OpenStudy (anonymous):

\[\frac{83}{84}\]

OpenStudy (anonymous):

the answer is what i put last... 83/84

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

83/84

OpenStudy (precal):

reduce your fraction what number can you divide the top and bottom?

OpenStudy (anonymous):

thank you all so much

OpenStudy (anonymous):

ok precal is it the number that can be divinded into the bottom number only with are the demonatores